18, 19, 21, 27, 20


1
Expert's answer
2022-05-02T01:46:02-0400
"mean=\mu=\dfrac{18+19+21+27+20}{5}=21"

"Var(X)=\sigma^2=\dfrac{9+4+0+36+1}{5}=10"

"\sigma=\sqrt{\sigma^2}=\sqrt{10}\approx3.1623"